OpenStudy (anonymous):

Solve the Equation Please √q-9 = 7

Parth (parthkohli):

square both sides q - 9 = 49 q = 58

OpenStudy (saifoo.khan):

Take square root on both sides, \[\sqrt{q-9} = 7\]\[q - 9 = 49\]\[q = 49 + 9\]\[q = 58\]

OpenStudy (anonymous):

Thanks you guys

Parth (parthkohli):

@nataliepaj : Give saifoo a medal, not me

OpenStudy (saifoo.khan):

It's okay. @ParthKohli . :)